### v1

@@ Abstract
The rapid growth of Internet of Things (IoT) devices produces massive, heterogeneous data streams, demanding scalable and efficient scheduling in cloud environments to meet latency, energy, and Quality-of-Service (QoS) requirements. Existing scheduling methods often lack adaptability to dynamic workloads and network variability inherent in IoT-cloud systems. This paper presents a novel hybrid scheduling algorithm combining deep Reinforcement Learning (RL) and Ant Colony Optimization (ACO) to address these challenges. The deep RL agent utilizes a model-free policy-gradient approach to learn adaptive task allocation policies responsive to real-time workload fluctuations and network states. Simultaneously, the ACO metaheuristic conducts a global combinatorial search to optimize resource distribution, mitigate congestion, and balance load across distributed cloud nodes. Extensive experiments on large-scale synthetic IoT datasets, reflecting diverse workloads and QoS constraints, demonstrate that the proposed method achieves up to 18.4% reduction in average response time, 12.7% improvement in resource utilization, and 9.3% decrease in energy consumption compared to leading heuristics and RL-only baselines. Moreover, the algorithm ensures strict Service Level Agreement (SLA) compliance through deadline-aware scheduling and dynamic prioritization. The results confirm the effectiveness of integrating model-free RL with swarm intelligence for scalable, energy-efficient IoT data scheduling, offering a promising approach for next-generation IoT-cloud platforms.
keywords:
[…]
1 Introduction
@@
The proliferation of Internet of Things (IoT) systems has led to an unprecedented surge in data generation, necessitating scalable and latency-aware computing infrastructures within cloud and grid computing environments [1, 2]. Traditional IoT cloud architectures are typically characterized by tightly coupled compute and storage components managed under centralized control. However, such designs struggle to accommodate the extreme heterogeneity, dynamic workload fluctuations, and stringent latency constraints inherent in modern large-scale, data-intensive IoT deployments [3, 4]. These limitations often manifest as scalability bottlenecks, network congestion, and high operational overhead, particularly in geographically distributed or resource-constrained settings.
We proposed to enhance data reliability and parallel processing capabilities of IoT clusters, which commonly adopt redundant storage strategies and replicate data blocks across multiple nodes. This approach enables local task execution; tasks are processed on nodes that host the required data, thus improving efficiency through reduced transmission latency. However, recent trends toward disaggregated cloud architectures, where compute and storage resources are decoupled and connected via ultra-high-speed interconnects [5, 6, 7], offer promising flexibility. These architectures support independent resource scaling, modular upgrades, and hardware heterogeneity, making them well-suited for complex IoT environments. However, they also introduce critical challenges related to non-local data execution and increased data movement across the network. Given the immense volume of real-time sensing data, network infrastructure often becomes a performance bottleneck due to bandwidth constraints and communication latency [8, 9].
The emerging computing and processing IoT frameworks increasingly adopt data-parallel processing paradigms that co-locate computation with data at the node level. Data block placement aligns task execution with data locality; these systems minimize network load and improve resource utilization [10, 11]. Convergent, realizing efficient task-to-data affinity in heterogeneous IoT clusters is highly non-trivial. Nodes differ significantly in computational capabilities, memory hierarchies, and I/O throughput data placement. Moreover, workloads are dynamic and unpredictable, with variations in task arrival rates, contention for shared resources, and energy constraints—further complicating scheduling decisions. Distributed file systems such as the Hadoop Distributed File System (HDFS) [12, 13], widely used in IoT infrastructures, replicate data blocks to improve fault tolerance and availability. However, this replication introduces a combinatorial explosion in task-to-node scheduling choices. Resource contention, naive heuristics, and static placement strategies can lead to degraded data locality and suboptimal system performance. Existing schedulers often rely on hierarchical locality models for execution on local or same-rack nodes. While effective in homogeneous environments, these models neglect node-specific capabilities in heterogeneous clusters, leading to severe load imbalance and scheduling inefficiencies [14, 15].
This paper proposes a novel challenge and a context-aware, reinforcement learning-driven scheduling framework tailored for heterogeneous IoT clusters. The proposed framework, Sensor Cloud Computing and Data Scheduling Optimization (SCC-DSO), jointly considers data placement topology, node-level resource heterogeneity, and workload dynamics to optimize task-to-node mapping [16]. The SCC-DSO formulates scheduling as a constrained min–max optimization problem aimed at minimizing makespan while maximizing data locality, subject to node capacity and network bandwidth constraints. A kernel-based regression model predicts task execution times using node and workload features, capturing complex non-linear relationships [17, 18]. SCC-DSO incorporates a metaheuristic strategy that fuses reinforcement learning with ant colony optimization to navigate the high-dimensional scheduling space. The proposed framework is designed for next-generation IoT applications that demand ultra-low latency, high reliability, and scalable resource utilization. These include mission-critical domains such as autonomous driving, industrial robotics, and environmental sensing. SCC-DSO aims to deliver robust performance across dynamic, distributed IoT-cloud infrastructures by enabling intelligent, adaptive scheduling in real-time [19, 20].
The primary contributions of this work are a way to go there: (i). We propose a novel integration of reinforcement learning (RL) and metaheuristic optimization to enable adaptive, heterogeneity-aware task allocation in IoT cloud environments, enhancing scheduling efficiency under dynamic workloads. (ii). We formulate a constrained min-max optimization model that jointly maximizes data locality and minimizes execution latency, addressing the challenges of resource heterogeneity and stringent latency requirements in IoT-cloud systems. (iii). We develop an RL-guided ant colony optimization (ACO) mechanism that facilitates adaptive task migration and predictive data prefetching, ensuring robust performance in highly dynamic and unpredictable workload scenarios. (iv). We validate the SCC-DSO framework through extensive experiments on a 100-node heterogeneous cluster, achieving up to 22.4% reduction in execution time, 93.1% data locality, and significant throughput improvements compared to state-of-the-art baselines, demonstrating its scalability, efficiency, and robustness for latency-sensitive IoT cloud applications [21, 22].
2 Related Work
The paper emphasized that in data placement optimization and computing execution, efficient data scheduling and resource optimization in heterogeneous Internet of Things (IoT) clusters remain central challenges in scalable edge–cloud computing. Moreover, various approaches have emerged to address these issues, particularly in the context of execution latency, energy efficiency, and resource heterogeneity. This section reviews key contributions across three primary dimensions: predictive task scheduling, data placement strategies, and energy-aware orchestration.
@@ 2 Related Work
Predictive Scheduling Models: One widely studied technique is the Reservation First-Fit with Feedback Distribution (RF-FD) model [23], which utilizes multiple linear regression to estimate job completion times based on historical execution patterns. While RF-FD performs well in semi-homogeneous clusters, its linear modeling lacks flexibility in handling node-level variability and non-linear workload dynamics. In contrast, RSYNC [24] employs a log-based synchronization mechanism to maintain consistency across fog nodes under fluctuating data streams. Although RSYNC is effective for sequential synchronization, it suffers from throughput degradation in high-load, heterogeneous environments. Autonomic frameworks such as those built on the MAPE-K loop [25], incorporating Monitoring, Analysis, Planning, Execution, and Knowledge, enhance adaptability through runtime feedback. However, their assumptions of static task profiles and homogeneous infrastructures limit applicability in dynamic IoT deployments. Machine learning-based solutions, such as the polynomial regression scheduler proposed in [26], offer non-linear adaptability by modeling task volume and node utilization. Nonetheless, their performance deteriorates under high-dimensional variance, highlighting the need for models with better generalization capabilities.
@@
Data Placement and Locality Optimization: Data replication and placement strategies are equally critical for minimizing network overhead and maximizing execution locality. The traditional rack-aware strategy is applied to the Hadoop Distributed File System (HDFS) [27, 28], replicating blocks locally, within the rack, and across racks to ensure fault tolerance and minimize latency. However, these uniform replication approaches often lead to load imbalance in heterogeneous environments, where node capabilities vary significantly [29, 30]. To address this, recent studies have explored dynamic data placement methods considering node-specific compute capacity, but few integrate these with intelligent scheduling to achieve end-to-end optimization. The SCC-DSO framework addresses this gap by aligning task assignment with predictive execution performance and real-time node capabilities.
@@
Energy-Efficient Scheduling and Virtualization: The scheduling of IoT systems that are resource-constrained is another critical area that requires efficient energy consumption. Constrained energy models [31], Dynamic Voltage and Frequency Scaling (DVFS) techniques, and queuing-based power optimization [32] aim to reduce energy consumption without violating service-level agreements (SLAs). Virtualization approaches, such as live virtual machine (VM) migration, contribute to energy efficiency by enabling workload consolidation and resource overcommitment. Despite these advances, existing predictive models based on hardware performance counters or statistical VM energy profiling offer limited scalability in distributed, real-time IoT systems due to their reliance on static or linear assumptions [33, 34].
@@
SCC-DSO Contribution: The proposed SCC-DSO framework advances the state of the art by holistically integrating kernel-based execution time prediction, reinforcement learning-driven adaptation, and metaheuristic optimization through Ant Colony Optimization (ACO). Its tri-layered architecture dynamically adapts to workload fluctuations, node heterogeneity, and data locality constraints in the IoT cluster. Unlike traditional schedulers that rely on reactive mechanisms or oversimplified models, SCC-DSO provides a robust, scalable, and energy-efficient foundation for intelligent task orchestration in complex edge–cloud IoT ecosystems [35, 36].
Figure 1: Hybrid IoT-cloud architecture integrates decentralized IPFS storage with centralized orchestration via MariaDB and MQTT, enabling scalable and reliable data management across the edge and cloud nodes.
@@
The proposed Figure 1 architecture presents a hybrid framework for heterogeneous IoT-cloud environments, integrating decentralized storage mechanisms with centralized orchestration to address key scalability and reliability challenges in modern sensor networks. Since, in the core principles, sensor nodes with edge devices capture multi-modal data streams ranging from audio to environmental to local brokers using the MQTT protocol. The broker interfaces with a traditional relational database (MariaDB) [37] and a distributed IPFS cluster, enabling dual-path data handling, and content identifiers (CIDs) generated at the edge are logged in the database for auditability. Consequently, data management is stored across a decentralized IPFS private swarm, ensuring tamper resistance, fault tolerance, and rapid retrieval. This hybrid model enhances data integrity, real-time responsiveness, and interoperability across diverse IoT nodes [38].
3 Proposed Methodology
@@
The Sensor Cloud Computing and Data Scheduling Optimization (SCC-DSO) framework addresses performance bottlenecks in heterogeneous IoT-edge clusters by integrating reinforcement learning (RL), ant colony optimization (ACO), and predictive modeling. This multi-stage algorithm dynamically schedules tasks based on data locality, node capability, and network conditions to meet stringent latency and throughput requirements [39]. The IoT cluster is modeled as a graph $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of computing nodes has heterogeneous processing capacity, memory, and I/O characteristics and $\mathcal{E}$ represents communication links. Input datasets are partitioned into fixed-size blocks (64 MB in HDFS), with replicas ensuring fault tolerance. Tasks are assigned to nodes hosting required data or fetching it with minimal overhead [40].
@@
$\sum_{i\in\mathcal{V}}x_{ij}$ $=1,\quad\forall j\in\mathcal{J},$ (2)
$\sum_{j\in\mathcal{J}}x_{ij}d_{j}$ $\leq C_{i}(t),\quad\forall i\in\mathcal{V},$ (3)
$x_{ij}$ $\in\{0,1\},\quad\forall i\in\mathcal{V},j\in\mathcal{J},$ (4)
@@
where $\mathbf{x}_{j}=[m_{j},\mathbf{z}_{i}]$ combines task data size $m_{j}$ and node features $\mathbf{z}_{i}$ (e.g., CPU speed, memory), $K(\mathbf{x}_{j},\mathbf{x}_{s})=\exp\left(-\frac{\|\mathbf{x}_{j}-\mathbf{x}_{s}\|^{2}}{2\sigma^{2}}\right)$ is a Gaussian radial basis function kernel, $\{(\mathbf{x}_{s},t_{s})\}_{s=1}^{S}$ are historical training data, $w_{s}$ are learned coefficients, $b$ is a bias term, and $\sigma\in[0.5,2.0]$ (Table 1) is the kernel bandwidth. The learning rate $\gamma\in[0.01,0.1]$ tunes the model’s convergence.
@@
3.1 Ant Colony Optimization (ACO) Scheduling Mechanism
SCC-DSO adapts ACO with predictive modeling and heterogeneity-aware pheromone updates to solve NP-hard task-to-node scheduling problems. The process has four phases. Initialization: A graph $G=(\mathcal{V},\mathcal{E})$ is initialized with pheromone trails $\tau_{ij}(0)=\tau_{0}\in[0.01,0.1]$ (Table 1), incorporating execution time predictions from Eq. (5). Solution Construction: Ants select task-node assignments using:
@@ 3.1 Ant Colony Optimization (ACO) Scheduling Mechanism
where $\eta_{ij}=1/T_{ij}$ is the heuristic desirability $\alpha\in[1.0,2.0]$ and $\beta\in[2.0,3.0]$ control pheromone and heuristic influence (Table 1), and $\mathcal{V}_{\text{eligible}}$ is the set of eligible nodes for task $j$.
@@
where $\rho\in[0.1,0.3]$ is the evaporation rate, $Q\in[100,500]$ is a constant, $K\in[10,20]$ is the number of ants, and $L^{(k)}$ is the makespan of the $k$-th ant’s schedule (Table 1). A minimum pheromone trail $\delta\in[10^{-4},10^{-2}]$ prevents stagnation.
@@
where $RQ_{i}>\varphi$ and $RQ_{i}-TS_{i}>\varphi$, with $\varphi\in[0.05,0.1]\cdot TS_{i}$, $r$ is the task’s resource demand, $B_{k}$ is the data block size, $\rho\in[0.1,0.3]$ is a scaling factor, and $V_{i}$ is the node’s processing rate (Table 1). The function $T(\cdot)$ was removed as it was undefined; the expression is simplified as a direct ratio. The optimal source node minimizes:
@@
Table 1: Parameters and Ranges for SCC-DSO
@@
Table 1 lists the configuration parameters for SCC-DSO. The pheromone influence $\alpha$ balances exploration and exploitation, while $\beta$ heuristic information is emphasized (e.g., execution time). The evaporation rate $\rho$ and initial pheromone $\tau_{0}$ regulate trail persistence. Parameters $K$, II, and $w_{1},w_{2},w_{3}$ balance computational overhead and performance. The prefetch threshold $\varphi$, learning rate $\gamma$, migration limit $\theta$, and kernel bandwidth $\sigma$ support predictive modeling and dynamic scheduling, ensuring scalability and adaptability in IoT-edge clusters.
3.2 Lightweight Hybrid RL-ACO Model
The Sensor Cloud Computing and Data Scheduling Optimization (SCC-DSO) framework introduces a lightweight hybrid Reinforcement Learning (RL) and Ant Colony Optimization (ACO) scheduler tailored for resource-constrained edge devices in IoT-edge clusters. This variant reduces computational overhead while maintaining high scheduling accuracy, making it suitable for latency-critical applications such as autonomous driving, industrial robotics, and smart surveillance systems.
3.2.1 Lightweight Scheduler Design

### v2

@@ Abstract
The rapid growth of Internet of Things (IoT) devices produces massive, heterogeneous data streams, demanding scalable and efficient scheduling in cloud environments to meet latency, energy, and Quality-of-Service (QoS) requirements. Existing scheduling methods often lack adaptability to dynamic workloads and network variability inherent in IoT-cloud systems. This paper presents a novel hybrid scheduling algorithm combining deep Reinforcement Learning (RL) and Ant Colony Optimization (ACO) to address these challenges. The deep RL agent utilizes a model-free policy-gradient approach to learn adaptive task allocation policies responsive to real-time workload fluctuations and network states. Simultaneously, the ACO metaheuristic conducts a global combinatorial search to optimize resource distribution, mitigate congestion, and balance load across distributed cloud nodes. Extensive experiments on large-scale synthetic IoT datasets, reflecting diverse workloads and QoS constraints, demonstrate that the proposed method achieves up to $18.4\%$ reduction in average response time, $12.7\%$ improvement in resource utilization, and $9.3\%$ decrease in energy consumption compared to leading heuristics and RL-only baselines. Moreover, the algorithm ensures strict Service Level Agreement (SLA) compliance through deadline-aware scheduling and dynamic prioritization. The results confirm the effectiveness of integrating model-free RL with swarm intelligence for scalable, energy-efficient IoT data scheduling, offering a promising approach for next-generation IoT-cloud platforms.
Index Terms:
Evolutionary Algorithms; Data Scheduling; Cloud Computing; Reinforcement Learning; Metaheuristics; Resource Optimization
I Introduction
@@
The rapid evolutionary computing of Internet of Things (IoT) systems has caused an unprecedented surge in data generation, demanding scalable and latency-aware computing infrastructures within cloud and grid environments [1, 2]. Conventional IoT cloud architectures rely on tightly coupled compute-storage units under centralized control, which struggle to handle extreme heterogeneity, dynamic workloads, and stringent latency constraints in large-scale, data-intensive IoT deployments [3, 4]. These limitations often result in scalability bottlenecks, network congestion, and increased operational overhead, especially across geographically distributed or resource-constrained settings. However, IoT clusters typically replicate data blocks across multiple nodes, enabling local task execution and reducing transmission latency to enhance data reliability and parallel processing. However, emerging disaggregated cloud architectures decouple compute and storage resources via ultra-high-speed interconnects, offering modularity and independent scaling [5, 6, 7]. Consequently, such designs exacerbate challenges related to non-local data execution and heightened network data movement, which frequently become bottlenecks given the massive real-time sensing data and bandwidth constraints [8, 9].
Data-parallel processing paradigms that co-locate computation with data have emerged to mitigate these issues by minimizing network load and optimizing resource utilization [10, 11]. Nonetheless, achieving efficient task-to-data affinity in heterogeneous IoT clusters remains challenging due to diverse node capabilities, dynamic workloads, and energy constraints. Distributed file systems such as HDFS [12, 13] replicate data blocks to improve fault tolerance, exponentially increasing scheduling computational complexity. Furthermore, existing locality-based schedulers often ignore node heterogeneity, causing load imbalance and suboptimal performance [14, 15]. This paper introduces Sensor Cloud Computing and Data Scheduling Optimization (SCC-DSO), a context-aware, reinforcement learning-driven scheduling framework tailored for heterogeneous IoT clusters [16]. SCC-DSO formulates scheduling as a constrained min–max optimization problem to minimize makespan and maximize data locality under resource and bandwidth constraints. It leverages a kernel-based regression model to predict execution times from node and workload features integration, reinforcement learning with ant colony optimization to efficiently explore the high-dimensional scheduling space for a high-performance computing [17, 18].
Designed for latency-critical IoT applications such as autonomous driving, smart manufacturing, and industrial robotics, the proposed SCC-DSO framework introduces a robust scheduling paradigm that adapts to system heterogeneity while maintaining low latency and energy efficiency [19, 20]. Unlike conventional schedulers that either overlook heterogeneity or rely solely on static heuristics, SCC-DSO integrates reinforcement learning (RL) with metaheuristic optimization to support adaptive task mapping in dynamic cluster environments. Extensive experiments on a 100-node heterogeneous testbed demonstrate up to 22.4% reduction in execution time, 93.1% task–data locality, and consistently higher throughput relative to state-of-the-art baselines [21, 22].
The primary contributions of this work are fourfold: (i) development of a novel RL–metaheuristic hybrid that enables heterogeneity-aware and latency-sensitive task allocation under dynamic workloads; (ii) formulation of a constrained min–max optimization model that balances locality and latency objectives while preserving scalability; (iii) introduction of an RL-guided ant colony optimization (ACO) mechanism to support proactive task migration and data prefetching in distributed settings; and (iv) comprehensive empirical validation across varying cluster sizes, replication factors, and straggler scenarios, confirming SCC-DSO’s superiority in efficiency, adaptability, and robustness for IoT–cloud environments. Together, these contributions position SCC-DSO as a scalable solution for next-generation latency-sensitive applications.
II Related Work
This work introduces efficient data scheduling and resource optimization challenges in scalable, heterogeneous IoT edge–cloud environments. However, prior research addresses these through predictive scheduling, data placement optimization, and energy-efficient orchestration, focusing on minimizing latency, balancing workloads, and reducing energy consumption.
@@ II Related Work
Predictive Scheduling Models: Early models like Reservation First-Fit with Feedback Distribution (RF-FD) [23] apply multiple linear regression to predict job completion based on historical data, performing well in semi-homogeneous clusters but lacking flexibility for non-linear workload variations. RSYNC [24] offers log-based synchronization across fog nodes but suffers throughput degradation under high load and heterogeneity. Autonomic frameworks based on MAPE-K loops [25] enhance adaptability through runtime feedback; however, they assume static task profiles and homogeneous infrastructures, limiting their use in dynamic IoT contexts. Machine learning approaches such as polynomial regression schedulers [26] improve non-linear modeling but face challenges with high-dimensional data variance, underscoring the need for more generalizable models.
@@
Data Placement and Locality Optimization: Data replication strategies like the rack-aware policy in Hadoop Distributed File System (HDFS) [27, 28] replicate data blocks locally and across racks to improve fault tolerance and reduce latency. These uniform approaches, however, often cause load imbalance in heterogeneous clusters with variable node capabilities [29, 30]. Recent work explores dynamic placement considering node compute capacity, yet integration with intelligent scheduling remains sparse. The SCC-DSO framework fills this gap by aligning task assignment with predictive execution time and node resource profiles, optimizing data locality and workload distribution in real-time optimization.
@@
Energy-Efficient Scheduling and Virtualization: Energy-aware techniques, including constrained energy models [31], Dynamic Voltage and Frequency Scaling (DVFS) [32], and queuing-based power optimization, aim to reduce energy use without SLA violations. Virtualization technologies such as live VM migration facilitate workload consolidation and energy savings. Despite these advances, many predictive models relying on hardware counters or VM energy profiles have limited scalability in distributed IoT due to static assumptions and linearity constraints [33, 34].
@@
SCC-DSO Contributions: Unlike prior work, SCC-DSO integrates kernel-based execution time prediction with reinforcement learning and metaheuristic Ant Colony Optimization (ACO) for dynamic, heterogeneity-aware scheduling. Its tri-layer architecture adapts to workload variability, node heterogeneity, and data locality constraints, enabling robust, scalable, and energy-efficient task orchestration in complex IoT-cloud systems [35, 36].
@@
Figure 1 illustrates a hybrid IoT-cloud architecture combining decentralized storage via IPFS with centralized orchestration. Edge sensor nodes generate multi-modal data streams processed locally and forwarded through MQTT brokers interfacing with MariaDB [37] and an IPFS private swarm, ensuring tamper resistance, fault tolerance, and rapid retrieval. This hybrid model enhances data integrity, responsiveness, and interoperability across heterogeneous IoT environments [38].
Figure 1: Hybrid IoT-cloud architecture combines IPFS, MariaDB, and MQTT for scalable, reliable edge-to-cloud data management.
III Proposed Evolutionary Algorithm
The proposed architecture 2 integrates a Hybrid Electro Search–Ant Colony Optimization (ES-ACO) algorithm, enabling efficient task scheduling in sensor cloud environments and delivering optimized performance for intelligent cloud-based irrigation control [69]. The diagram shows that cloud users submit tasks managed by cloud brokers and a task manager before forwarding them to the hybrid ES-ACO scheduler. This scheduler optimizes task allocation across virtual machines (VMs) hosted within a cloud data center and mapped to physical sensor nodes via resource managers’ integration of Wireless Sensor Networks (WSNs) data collection from a distributed computing environment. The multi-objective scheduler aims to minimize energy consumption, make span, and execution cost while maximizing throughput and reducing task rejection ratio. This approach ensures intelligent irrigation control, resource-efficient, and sustainable agricultural practices.
@@ III Proposed Evolutionary Algorithm
The Sensor Cloud Computing and Data Scheduling Optimization (SCC-DSO) framework addresses performance bottlenecks in heterogeneous IoT-edge clusters by integrating reinforcement learning (RL), ant colony optimization (ACO), and predictive modeling. This multi-stage algorithm dynamically schedules tasks based on data locality, node capability, and network conditions to meet stringent latency and throughput requirements [39]. The IoT cluster is modeled as a graph $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of computing nodes that have heterogeneous processing capacity, memory, and I/O characteristics, and $\mathcal{E}$ represents communication links. Input datasets are partitioned into fixed-size blocks (64 MB in HDFS), with replicas ensuring fault tolerance. Tasks are assigned to nodes hosting required data or fetching it with minimal overhead [40].
Figure 2: Hybrid ES-ACO Algorithm for Task Scheduling in a Sensor Cloud for Smart Irrigation
@@
$\sum_{i\in\mathcal{V}}x_{ij}$ $=1,\quad\forall j\in\mathcal{J},$
$\sum_{j\in\mathcal{J}}x_{ij}d_{j}$ $\leq C_{i}(t),\quad\forall i\in\mathcal{V},$
$x_{ij}$ $\in\{0,1\},\quad\forall i\in\mathcal{V},j\in\mathcal{J}.$
@@
where $\mathbf{x}_{j}=[m_{j},\mathbf{z}_{i}]$ combines task data size $m_{j}$ and node features $\mathbf{z}_{i}$ (e.g., CPU speed, memory), $K(\mathbf{x}_{j},\mathbf{x}_{s})=\exp\left(-\frac{\|\mathbf{x}_{j}-\mathbf{x}_{s}\|^{2}}{2\sigma^{2}}\right)$ is a Gaussian radial basis function kernel, $\{(\mathbf{x}_{s},t_{s})\}_{s=1}^{S}$ are historical training data, $w_{s}$ are learned coefficients, $b$ is a bias term, and $\sigma\in[0.5,2.0]$ (Table I) is the kernel bandwidth. The learning rate $\gamma\in[0.01,0.1]$ tunes the model’s convergence.
@@
III-A Ant Colony Optimization (ACO) Scheduling Mechanism
SCC-DSO adapts ACO with predictive modeling and heterogeneity-aware pheromone updates to solve NP-hard task-to-node scheduling problems. The process has four phases. Initialization: A graph $G=(\mathcal{V},\mathcal{E})$ is initialized with pheromone trails $\tau_{ij}(0)=\tau_{0}\in[0.01,0.1]$ (Table I), incorporating execution time predictions from Eq. (2). Solution Construction: Ants select task-node assignments using:
@@ III-A Ant Colony Optimization (ACO) Scheduling Mechanism
where $\eta_{ij}=1/T_{ij}$ is the heuristic desirability $\alpha\in[1.0,2.0]$ and $\beta\in[2.0,3.0]$ control pheromone and heuristic influence (Table I), and $\mathcal{V}_{\text{eligible}}$ is the set of eligible nodes for task $j$.
@@
where $\rho\in[0.1,0.3]$ is the evaporation rate, $Q\in[100,500]$ is a constant, $K\in[10,20]$ is the number of ants, and $L^{(k)}$ is the makespan of the $k$-th ant’s schedule (Table I). A minimum pheromone trail $\delta\in[10^{-4},10^{-2}]$ prevents stagnation.
@@
where $RQ_{i}>\varphi$ and $RQ_{i}-TS_{i}>\varphi$, with $\varphi\in[0.05,0.1]\cdot TS_{i}$, $r$ is the task’s resource demand, $B_{k}$ is the data block size, $\rho\in[0.1,0.3]$ is a scaling factor, and $V_{i}$ is the node’s processing rate (Table I). The function $T(\cdot)$ was removed as it was undefined; the expression is simplified as a direct ratio. The optimal source node minimizes:
@@
TABLE I: Parameters and Ranges for SCC-DSO
@@
Table I lists the configuration parameters for SCC-DSO. The pheromone influence $\alpha$ balances exploration and exploitation, while $\beta$ heuristic information is emphasized (e.g., execution time). The evaporation rate $\rho$ and initial pheromone $\tau_{0}$ regulate trail persistence. Parameters $K$, II, and $w_{1},w_{2},w_{3}$ balance computational overhead and performance. The prefetch threshold $\varphi$, learning rate $\gamma$, migration limit $\theta$, and kernel bandwidth $\sigma$ support predictive modeling and dynamic scheduling, ensuring scalability and adaptability in IoT-edge clusters.
III-B Lightweight Hybrid RL-ACO Methods
The Sensor Cloud Computing and Data Scheduling Optimization (SCC-DSO) framework incorporates a lightweight hybrid scheduler that combines Reinforcement Learning (RL) with Ant Colony Optimization (ACO), specifically designed for resource-constrained edge devices in IoT–edge clusters. This approach minimizes computational overhead while preserving high scheduling accuracy, making it well-suited for latency-sensitive applications such as autonomous driving, industrial robotics, and intelligent surveillance systems. RL is employed to learn adaptive task-node selection policies based on dynamic network conditions, while ACO refines these selections through heuristic-guided pheromone updates to achieve near-optimal scheduling decisions 1.
Computational Complexity: Let $|T|$ be the number of tasks, $|N|$ the number of nodes, and $I$ the number of ACO iterations. The RL policy inference operates in $\mathcal{O}(|T||N|)$ per scheduling round, while the ACO refinement contributes $\mathcal{O}(I|T||N|)$ due to pheromone and heuristic updates. Thus, the overall complexity is:
$\mathcal{O}(|T||N|(1+I))$
Given that $I\ll|T|$ in our lightweight configuration, the scheduler maintains near-linear scalability concerning the task volume.
Input: Task queue $T$, node set NN, resource capacities $R$
Output: Optimized task-node assignment
Initialize RL policy $\pi_{\theta}$ and pheromone matrix $\tau$;
while Tasks remain in $T$ do
  Extract state features from the current cluster load;
  Select candidate task-node pairs using $\pi_{\theta}$;
  Apply ACO refinement:
  1. 1.
    Compute heuristic desirability $\eta_{ij}$ for each task-node pair
  2. 2.
    Update pheromone trails $\tau_{ij}$ based on solution quality
  3. 3.
    Select path with maximum $\tau_{ij}\cdot\eta_{ij}$
Execute selected tasks locally where possible;
  Update $\pi_{\theta}$ using observed latency and resource utilization;
return Optimized assignment plan;
Algorithm 1 Lightweight Hybrid RL-ACO Scheduler
Latency Model: The end-to-end scheduling latency $L_{total}$ is modeled as:
$L_{total}=L_{comp}+L_{trans}+L_{queue}$
where $L_{comp}$ denotes computation delay for policy inference and ACO updates, $L_{trans}$ is the transmission delay between nodes, and $L_{queue}$ is the queuing delay before task execution. By enabling local execution and reducing $L_{trans}$, the proposed method achieves substantial latency savings:
$L_{total}^{\text{Hybrid}}<L_{total}^{\text{Baseline}}$
under identical workload and network conditions.
III-B1 Lightweight Scheduler Design